### v1

# Biologically-Inspired Technologies: Integrating Brain-Computer Interface and Neuromorphic Computing for Human Digital Twins

Abstract
The integration of the Metaverse into a human-centric ecosystem has intensified the need for sophisticated Human Digital Twins (HDTs) that are driven by the multifaceted human data. However, the effective construction of HDTs faces significant challenges due to the heterogeneity of data collection devices, the high energy demands associated with processing intricate data, and concerns over the privacy of sensitive information. This work introduces a novel biologically-inspired (bio-inspired) HDT framework that leverages Brain-Computer Interface (BCI) sensor technology to capture brain signals as the data source for constructing HDT. By collecting and analyzing these signals, the framework not only minimizes device heterogeneity and enhances data collection efficiency, but also provides richer and more nuanced physiological and psychological data for constructing personalized HDTs. To this end, we further propose a bio-inspired neuromorphic computing learning model based on the Spiking Neural Network (SNN). This model utilizes discrete neural spikes to emulate the way of human brain processes information, thereby enhancing the system’s ability to process data effectively while reducing energy consumption. Additionally, we integrate a Federated Learning (FL) strategy within the model to strengthen data privacy. We then conduct a case study to demonstrate the performance of our proposed twofold bio-inspired scheme. Finally, we present several challenges and promising directions for future research of HDTs driven by bio-inspired technologies.
Index Terms:
Human Digital Twin, Human-Centric Metaverse, Brain-Computer Interface, Spiking Neural Networks.
I Introduction
The Metaverse, once a concept confined to science fiction, has become a tangible reality. This shared virtual environment, created by the convergence of digitally enhanced physical reality, augmented reality (AR), and ubiquitous Internet access, allows users to engage in interactive, work-related, and social activities within three-dimensional spaces that transcend physical and geographic boundaries. With the advent of Beyond 5G (B5G) and 6G technologies, the Metaverse is evolving into a more human-centric space, i.e., human-centric Metaverse. This new paradigm aims not only to replicate but also to enhance human interactions through sophisticated digital extensions, ensuring that personal and collective engagements within these virtual spaces are as nuanced and meaningful as those in the physical world. To this end, the human digital twins (HDTs) has been proposed to thoroughly bridge the virtual and physical worlds [1].
Different from the virtual representation of users in the Metaverse (i.e., avatars), the HDT is an advanced system that extends beyond mirroring physical appearances. This system integrates and analyzes comprehensive data types of user, including behavioral, psychological, physiological, and observational information, enabling the HDT to adapt dynamically not only to user interactions and immersion within the Metaverse but also to provide real-time monitoring of human behavior and states. By leveraging each user’s unique and real-time data, HDTs efficiently capture real-time feedback and delivers personalized services, offering tailored solutions that meet the needs of both individual users and Metaverse Service Providers (MSPs). Therefore, this system not only significant elevates the Quality of Experience (QoE) perceived by the users but also assists in optimizing the Quality of Service (QoS) provided by the MSPs, ensuring efficient system performance. These multiple benefits highlight the integral role of HDTs in enhancing the overall functionality and sustainability of the human-centric Metaverse.
To construct an effective HDT, users must deploy multiple Internet-of-Things (IoT) devices (e.g., smartwatches, smartphones, and body sensors) to continuously collect real-time data on their physiological and behavioral states. This data is then transmitted to an MSP for comprehensive analysis, enabling the creation of highly personalized HDTs. However, this process presents significant challenges in two main stages respectively, i.e., data collection and data processing. Regarding data collection, the IoT devices used in this process are typically produced by different manufacturers, each operating on distinct software platforms and employing. This diversity complicates the efficient integration and processing of data. Varied operational protocols and data formats can lead to inconsistencies and gaps in the collected data, thereby reducing the efficacy of data collection. Additionally, the sensors in these devices are typically designed to detect only specific physiological signals, such as heart rate and blood oxygen levels. However, this may be insufficient in the dynamic real-world environments where human behaviors and states, such as psychological states like emotions and physiological states like fatigue, are complex and variable. Such sensors have limitations in capturing nuanced psychological or emotional states, which are essential for the immersive human-centric ecosystem.
On the other hand, the heterogeneity of IoT devices presents significant challenges for MSPs in efficiently processing data. MSPs must employ sophisticated algorithms capable of handling the diverse types of data generated by these devices. This necessity not only complicates the data integration process, requiring robust solutions to manage different operational protocols and data formats, but also impacts the system’s efficiency. Such complexities can diminish the accuracy and responsiveness of HDT, potentially affecting the immersive user experience. Furthermore, the vast amount of real-time data collected by numerous IoT devices, combined with their rapidly expanding user base, further poses significant pressure on data collection and processing systems while also heightening concerns about energy consumption and data security. Specifically, the intensive data handling required not only consumes considerable energy but also imposes strict constraints on the operational efficiency of these systems, thereby limiting the range of practical applications available to users. Although deep learning model are capable of managing these large data volumes by enhancing processing efficiency, their substantial energy requirements and the inherent risks of data breaches pose serious sustainability challenges.
Given these pivotal challenges identified in the development of HDTs, this work aims to answer several key research questions essential for advancing the field:
  1. Q1)
    How to accurately capture human activities data for constructing a HDT that dynamically adapts to users’ psychological and physiological changes, while reducing reliance on heterogeneous devices, thereby improving both the QoS for MSPs and the QoE for users?
  2. Q2)
    Given the need for continuous updating of HDT to adapt to real-time changes of user, what strategies can be developed to reduce the energy consumption involved in these data-intensive operations?
  3. Q3)
    Given the sensitivity of the individual data involved, how can security measures be enhanced during the continuous updates of HDT to protect against data breaches and efficiently manage vast volumes of sensitive data?
To address these challenges, we revisit current HDT frameworks from the perspective of human-centric Metaverse requirements. We then introduce a novel HDT framework and data processing method driven by biologically inspired (Bio-inspired) technologies. Our contributions are summarized as follows:
  • We introduce a novel Bio-inspired HDT (Bio-HDT) framework that utilizes Brain-Computer Interface (BCI) sensor technology to collect human brain signals as the primary data source for constructing HDTs. This framework harnesses brain signals to capture real-time behavioral, psychological, physiological, and observational information of individuals, thereby providing richer data and a holistic view for constructing immersive and personalized HDTs. This approach not only enhances the precision and fidelity of the HDTs but also significantly reduces the reliance on heterogeneous IoT devices, simplifying both data collection and processing. Consequently, it ensures high QoS provided by MSPs and QoE perceived by users (For Q1).
  • We propose a Bio-inspired neuromorphic computing learning model based on the Spiking Neural Networks (SNNs) for efficiently processing extracted brain signals. Unlike traditional Artificial Neural Networks (ANNs) that process information continuously and require constant computational activity, SNNs utilize discrete, brain-like spikes to transmit information between neurons. This method more accurately reflects the natural communication processes of brain neurons, thereby enhancing data processing efficiency and reducing the energy consumption specifically involved in the real-time updating of HDTs (For Q2).
  • We introduce an SNN-driven FL model that reduces raw data transmission to address privacy challenge (For Q3). Additionally, we conduct a case study to validate the performance of our proposed Bio-HDT framework and bio-inspired learning model.
To the best of our knowledge, this is the first work to explore the potential of Bio-inspired technologies in tackling the challenges faced by HDTs and the broader human-centric Metaverse.
II Human Digital Twin
In this section, we introduce the human-centric Metaverse and review the current HDT frameworks and their applications. Then, we highlight the key challenges that must be addressed for them to realize their full potential within the human-centric Metaverse.
II-A Human-centric Metaverse and Human Digital Twin Framework
As the next generation of the Internet, the human-centric Metaverse extends beyond providing a virtual space for gaming. It is envisioned as a new paradigm that leverages users’ behavioral, psychological, physiological, and observational data to bridge the virtual and physical worlds, thereby profoundly transforming human lifestyles. HDTs, similar to the critical role of digital twins (DT) in manufacturing [2], play a vital role in improving the human-centric Metaverse. For instance, they allow simulation and prediction of physiological and psychological states of humans, facilitating personalized medical treatments [1] and health management programs [3]. Besides, HDTs are also able to assist in generating intelligent avatars, creating highly customized environments that mimic real-world interactions and scenarios, thus offering unparalleled opportunities for users.
The current HDT framework operates cyclically by executing the following steps: (1) Sensors embedded in IoT devices continuously collect user real-time behavioral data; (2) The data are then transmitted to the MSP; (3) The MSP processes and analyzes the data using advanced software and algorithms. After analyzing these data, MSPs are able to generate the user’s HDT. Currently, this framework has been utilized primarily in healthcare and gaming. For example, HDTs are integrated with wearable devices such as smartwatches, fitness trackers, and health sensors to continuously monitor vital factors like heart rate, blood pressure, and glucose levels. This data can be synchronized with medical records and utilized to generate real-time insights into an individual’s health, enabling early diagnosis of potential issues and the optimization of personalized healthcare strategies [1, 3, 4].
II-B Challenges of the Current Human Digital Twins
While the concept of HDTs offers immense advantages, it is still in its early stages. Many ongoing multifaceted research challenges need to be addressed to fully realize its potential.
II-B1 Devices and Data Sources (Q1)
One of the significant challenges facing HDTs is the heterogeneity of the IoT devices used to collect human data. These IoT devices such as smartwatches, body sensors, smartphones, and fitness trackers are critical for collecting continuous streams of physiological data. However, these devices often produced by different manufacturers and have varying technical specifications. They operate on diverse hardware platforms, utilize different communication protocols, and employ distinct software architectures, creating significant barriers to seamless data integration and real-time processing. This heterogeneity complicates the creation of accurate, responsive HDTs, as each device requires specific processing capabilities and tailored integration efforts. Moreover, the need for real-time synchronization of data from multiple sources introduces additional technical complexities, including latency, bandwidth constraints, energy, and data loss, all of which degrade system performance. On the other hand, inconsistent data formats and interoperability issues increase the difficulty for MSPs in developing standardized frameworks for HDT implementation, thereby limiting the scalability and personalization potential of HDTs.
Furthermore, while these existing HDT frameworks can effectively replicate certain physiological functions (e.g., heart rate and blood oxygen), they are still far from achieving the comprehensive and dynamic representation of human bodies and minds that are necessary for fully immersive and personalized experiences in the human-centric Metaverse. The integration of advanced data sources, such as neurological and psychological inputs, remains nascent. This underscores a significant gap in the potential for HDTs to evolve into fully interactive and responsive entities within the human-centric Metaverse.
II-B2 Energy Consumption (Q2)
To efficiently analyze these collected data, advanced models and algorithms based on deep learning have been deployed. However, they are primarily driven by ANNs, which are resource-intensive and require substantial computational power for both training and inference phases [5]. This intense energy demand presents significant challenges for these IoT devices with limited energy. First, the high power requirement for real-time data processing and analysis can lead to frequent battery depletion, disrupting continuous monitoring and data collection essential for maintaining accurate and dynamic HDTs. Moreover, the strain on device endurance due to persistent high-load operations can degrade device performance over time, increasing maintenance costs and potentially compromising the quality of user interaction within the Metaverse.
II-B3 Data Security and Privacy (Q3)
HDTs are continuously updated with real-time data from IoT devices, introducing inherent risks of data breaches. The complexity of HDTs and the increase in data access points, such as sensors, wearable devices, and edge servers, expand potential cyber threats. Ensuring data security and privacy while maintaining HDT efficiency and responsiveness is a substantial technical challenge.
Based on the analysis above, it is evident that while HDTs have established a solid foundation for their role in certain applications, significant challenges persist. To address these challenges, we propose a novel HDT framework in the next section.
III Brain-Computer Interface-Enhanced Human Digital Twin
In this section, we introduce BCI technology and its current applications across various domains. We then propose a novel Bio-HDT framework that leverages advanced BCI technology, focusing on the potential for improved customization, efficiency, and interaction within the human-centric Metaverse. Finally, we discuss the unique challenges involved in deploying BCIs within HDT frameworks.
Figure 1: An illustration of the proposed framework for the biologically-inspired human digital twin (Bio-HDT). This process begins with the extraction of brain signals by a Brain-Computer Interface-driven Head-mounted Display (BCI-driven HMD). The signals are analyzed using the Spiking Neural Network to identify human behavior and states such as psychological state like stress and physiological state like fatigue, which are then mapped into the HDT. This framework enables MSPs to deliver personalized services tailored to the user’s real-time activities within a human-centric Metaverse.
III-A Brain Signal and Brain-Computer Interface
Unlike previous works [1, 3, 4], this work leverages brain signals as a novel data source for constructing HDTs. Specifically, brain signals offer a multifaceted view into human nature, crucial for modeling the human activities. They contain critical behavioral, psychological, physiological, and observational information, enabling the decoding of subconscious and conscious responses to environmental stimuli. These insights serve as windows into the internal processes that govern human actions, perceptions, and interactions, presenting their significant potential for constructing HDTs. By harnessing these signals, we can develop a human-centric Metaverse where each user’s HDT is authentically driven by their unique brain activities, ensuring a more immersive and personalized virtual experience.
To this end, BCI technology that connects the brain signals and physical devices has been considered. As a direct communication pathway between the human brain and externals machines, devices, and system, BCI operates by monitoring brain activities and collecting brain signals through sensors attached to the human scalp (e.g., non-invasive BCI [6]), processing the data to detect specific patterns related to intended actions, and translating these patterns into actionable commands. This allows users to control devices or perform tasks without physical movement [7]. It has been successfully applied in various domains, demonstrating its broad potential and utility.
  • Medical: BCIs provide individuals with spinal cord injuries or neurodegenerative diseases greater autonomy by enabling control over prosthetic limbs, communication systems, and other assistive technologies. Additionally, BCIs are used in clinical settings to assess brain injuries and diagnose brain death, aiding in critical decision-making processes in healthcare [7].
  • Mental Health: BCIs are transforming the diagnosis and treatment of mental health disorders, including conditions like epilepsy, strokes, and sleep disorders. Their ability to detect subtle neurological signals enables them invaluable in managing neurodegenerative diseases such as Parkinson’s [8] and Alzheimer’s [9].
  • Entertainment: BCIs are revolutionizing gaming and immersive experiences by enabling users to control avatars and interact with virtual environments using only their thoughts. This technology significantly enhances user immersion and interaction, setting the stage for more intuitive and responsive VR systems [10, 6, 11].
Unlike standard physical sensors technology that detect straightforward parameters such as temperature, pressure, or light, the human brain processes a rich and complex array of signals related to emotions, cognition, and behaviors. This comprehensive capability enables BCIs to access a spectrum of human states data that conventional sensors typically cannot, e.g., physical, mental, emotional, and behavioral states. Therefore, BCI offers the significant potential to create highly sophisticated and biologically accurate digital representations within the Metaverse, capturing subtleties of the human states that are beyond the reach of physical sensors.
III-B Biologically-inspired Human Digital Twins
As analyzed above, brain signals encompass underlying information about human activities. These activities provide a holistic view of an individual’s real-time states across multiple dimensions, which is particularly useful in contexts such as healthcare, personalized technology interactions, and immersive Metaverse application. By capturing and analyzing these various human behavior and states information, BCI technology assists in creating more personalized and responsive systems that adapt to individual needs and changes in their condition. Therefore, we propose the Bio-HDT framework that utilizes brain signals as the data source for conducting HDT and incorporates BCI technology to efficiently capture these signals. As illustrated in Fig. 1, the BCI-driven Head-mounted Display (HMD) system within the Bio-HDT framework continuously captures brain signals through embedded BCI sensors, which are then analyzed using a neuromorphic computing learning model that mimics brain function through SNNs. After real-time extraction and analysis of these brain signals, the Bio-HDT is constructed and synchronized with the corresponding human states in the physical world. Utilizing this real-time Bio-HDT, MSPs can identify and even predict fluctuations in a user’s psychology and physical intentions. Consequently, this framework is able MSPs to generate personalized services for the users by monitoring their brain signals (For Q1).
Specifically, the Bio-HDT operates as a two-layer bio-inspired scheme. The first layer bio-inspired captures brain signals, which reflect the user’s cognitive, emotional, and behavioral states, providing an authentic representation of the brain’s biological processes. These brain signals are continuously monitored and collected through embedded BCI sensors in the HMD system, ensuring real-time feedback from the user. The second layer bio-inspired processes these brain signals using SNNs, a neuromorphic computing scheme that mimics the way neurons communicate in the brain. Different from traditional neural networks, SNNs process data in a manner that mirrors the brain’s natural signaling mechanisms, interpreting neural patterns through the timing and frequency of spikes, which is detailed in Section IV. This allows MSPs to accurately extract, analyze, and respond to users’ brain signals, generating personalized content in real time.
By efficiently combining the BCI technology, the Bio-HDT framework offers several advantages. First, it minimizes the reliance on multiple IoT devices and external sensors traditionally required to monitor a user’s physical state. Conventional HDT systems rely on complex sensor networks, resulting in hardware demands, synchronization challenges, and increased energy consumption. By directly leveraging brain signals through the BCI, the Bio-HDT framework gathers critical multifaceted data (e.g., cognition, emotion, and behavior) without relying on an array of external sensors, thereby eliminating inefficiencies, reducing the investment in devices, and lowering computational load and data breach risk. Moreover, these brain signals are extracted unconsciously, without requiring explicit physical actions, preserving the user’s immersive experience and providing a more nuanced representation of their state. This facilitates more natural and immersive interactions within the Metaverse.
Furthermore, the Bio-HDT framework establishes a bi-directional feedback loop between the user and the Metaverse. As users interact with the virtual world, their brain signals influence their HDT’s behavior and responses. Conversely, the virtual environment provides stimuli or feedback that affects the user’s cognitive and emotional state, creating a continuous interaction between the physical and virtual worlds. As a result, MSPs can provide personalized services based on an individual’s unique Bio-HDT. For examples, MSPs can dynamically adapt the appearance, actions, or behavior of an avatar in Metaverse based on the user’s real-time physiological or psychological states, i,e., Bio-HDT. Besides, MSPs can also adjust the rendering schemes of VR content such as simplifying graphics and reducing interactive elements within the Metaverse to lessen cognitive load when detecting signs of stress and fatigue in the user.
Therefore, the Bio-HDT framework not only addresses the challenge of customizing HDTs for individual users (for Q1) but also minimizes energy consumption (for Q2) and reduces the risk of data breaches (for Q3) by eliminating the need for multiple data collection devices. Additionally, the real-time interaction between brain signals and the decision-making systems enhances the accuracy, responsiveness, and immersiveness of HDTs. This new paradigm brings users closer to a fully realized human-centric Metaverse.
III-C Concerns
Although Bio-HDT framework that utilizes brain signals and BCI technology offers rich information for updating HDTs and unprecedentedly supporting the development of the human-centric Metaverse, several concerns must be addressed to fully realize its potential.
III-C1 Efficient Processing of Brain Signals
Brain signals are inherently temporal and dynamic, necessitating that deep learning models process large volumes of data to decipher fluctuating patterns and rhythms of brain activity. This complexity significantly complicates real-time interpretation. Traditional ANNs, which typically rely on static data inputs, often struggle to manage these time-dependent fluctuations effectively. This inefficiency can lead to decreased accuracy in analysis and misinterpretations of the user’s states, potentially resulting in services that do not meet user expectations.
III-C2 Energy Efficiency (Q2)
Although the BCI-driven HMD alleviates the complexity of managing batteries for multiple sensors, processing real-time brain signals and constantly updating HDTs still requires significant computational power and energy resources. Current neural network architectures is energy-intensive, limiting their scalability and feasibility, particularly in large-scale, real-time applications within the Metaverse. Ensuring energy efficiency is critical to the sustainable operation of the Bio-HDT framework.
III-C3 Data Privacy and Security (Q3)
Given the sensitive and personal nature of brain signals, ensuring their secure transmission and processing is crucial. Transmitting these signals directly over wireless networks presents significant challenges due to the susceptibility of these channels to noise and interference, which can greatly degrade the quality and integrity of the data.
Therefore, it is an urgent need to design efficient, secure, and scalable approaches for processing brain data in real-time. In the following section, we introduce an SNN-driven FL model designed to enhance Bio-HDT framework. This model offers a robust solution that not only enhances the efficacy of brain data processing but also minimizes energy consumption and bolsters privacy protections.
IV Spiking Neural Network Enhanced Bio-HDT
Figure 2: The proposed network architecture based on SNNs. (a) Brain signal extraction is performed by HMD equipped with EEG electrodes to capture brain activity, (b) The captured brain signal is then input into a spike encoder, which converts the signal into spikes for further processing, (c) and (d) A brain signal identifier processes these firing spikes to accomplish brain signals identification.
In this section, we detail the SNN and introduce the SNN-driven FL to enhance the Bio-HDT framework. We then present a case study to support our proposed bio-inspired scheme.
IV-A Spiking Neural Network
SNN is a bio-inspired neuromorphic computing model that closely mimics the way of human brain processes information. Different from traditional neural networks (i.e., ANNs) that process information continuously and require constant computational activity, SNNs leverage discrete spikes of electrical activity to transmit information between neurons, more accurately reflecting the communication process of biological neurons [12, 13].
In the context of the Bio-HDT framework and the human-centric Metaverse, SNNs offer significant advantages due to their ability to process information based on the timing and frequency of spikes, closely mirroring the brain’s natural information processing mechanisms. These capabilities enables SNNs particularly adept at interpreting rapidly fluctuating brain signals and recognizing dynamic, time-based patterns that traditional models often overlook. As illustrated in the step (2) of Fig. 1, extracted brain signals are first fed into a brain-like SNN model. This model learns to process and efficiently encode brain activity into multiple spikes, enhancing the real-time responsiveness and accuracy in user states analysis and personalized service delivery. Importantly, the event-driven nature of SNNs, i.e., activating only during specific spike events, significantly reduces energy consumption and computational overhead, enabling SNNs ideally suited for power-sensitive devices such as HMDs and sensors within the Bio-HDT framework (for Q2). More technical details of SNN can be found in  [13].
IV-B SNN-Driven Federated Learning
MSPs traditionally collect users’ data to train deep learning models for states analysis. However, transmitting raw brain signals over wireless systems poses significant privacy concerns due to the sensitive and personal nature of this data. To address this challenge, we introduce the SNN-driven FL model. FL is a distributed machine learning approach where models are trained locally on user devices, and only the necessary updates are shared with a central server. This method eliminates the requirement to transmit raw data over the network, thereby enabling user privacy preservation while still enabling the collaborative training of models [14].
Specifically, SNN-driven FL can be deployed on local devices, i.e., BCI-driven HMDs. Instead of transmitting raw data to MSPs, each device trains its own SNN model using an FL framework. Note that model training occurs dynamically as users interact with the Metaverse, utilizing idle times or periods of low interaction to process data and refine the model. This strategy minimizes disruptions and maximizes efficiency during periods of active use. Once the local models are sufficiently trained, only the model updates (i.e., gradients) are transmitted to the MSP. There, they are aggregated to improve model performance without exposing the original brain data [14]. By retaining raw brain signals on local devices, the system protects users’ cognitive, emotional, and behavioral data from potential security threats during transmission. This architecture not only ensures a higher level of data privacy but also addresses critical concerns about the vulnerability of sensitive brain data (for Q3).
IV-C Case Study
By introducing the SNN-driven FL for Bio-HDT, the potential challenges aforementioned are efficiently addressed. We demonstrate the efficacy of our framework through comparison simulations in this section.
As aforementioned, the Bio-HDT framework is primarily designed to enhance three key factors: personalization, energy efficiency, and privacy preservation. In the human-centric Metaverse, personalization relies heavily on the accurate interpretation of human’s behavior intention and states, ensuring that user interactions are not only immersive but also highly responsive. For MSPs, the ability to accurately identify and interpret these signals is crucial. This capability is fundamental to effectively analyzing user requirements and creating personalized services that meet individual needs. The precision in processing these signals directly influences the quality, relevance, and timeliness of the content delivered. Consequently, it impacts both the QoS provided by the MSPs and the QoE perceived by the users. Additionally, energy efficiency is crucial for enabling sustainable and scalable operations in real-time environments, ensuring that the system can operate continuously without excessive resource consumption. Therefore, we evaluate our scheme’s performance based on the identification accuracy of brain signals and the energy consumption of neural network operations. Note that privacy is not evaluated in this case study because our SNN-driven FL framework effectively addresses privacy concerns by processing data locally. The settings of simulation are as follows:
IV-C1 Bio-HDT setting
[…]
IV-C2 SNN architecture
We conduct the learning model through SNN as shown in Fig. 2. This model comprises a spiking encoder network and a classifier network. The spiking encoder is designed to mimic neuronal communication, extracting features from brain signals and converting them into firing spikes at various time steps. The identifier network processes these spikes to carry out identification tasks specifically tailored to brain signals. Specifically, efficient identification of brain signals enables MSPs to understand users’ expectations, thus allowing them to provide personalized services.
IV-C3 FL training and baselines
We utilize the FedAvg algorithm [14] to conduct the FL framework. For the learning hyper-parameters, the learning rate, batch size and local epoch are set as $\left\{0.01,64,1\right\}$. Moreover, we conduct performance comparisons between our proposed method, denoted as FL4BCI-SNN, and two baseline methods: (i) FL4BCI-CNN and (ii) FL4BCI-LSTM. Specifically, the architecture of FL4BCI-CNN is designed to match that of FL4BCI-SNN, while the FL4BCI-LSTM network consists of two LSTM layers followed by a single linear layer.
Figure 3: Performance on various methods. (a) Energy consumption of various methods and (b) Identification accuracy vs. training round for various methods.
[…]
As aforementioned, both identification accuracy and energy consumption play a crucial role in determining the QoS provided by MSPs and the QoE experienced by users. As a result, we introduce a new metric Weighted System Performance (WSP) that combines identification accuracy and energy consumption. By integrating these factors into this metric and weighting them equally, we can ensure a balanced evaluation of user satisfaction across various system configurations. This metric allows for a comprehensive analysis of the trade-offs between identification accuracy and energy efficiency in service delivery, providing a holistic view of overall system performance.
Figure 4: The achievable WSP of various methods.
As illustrated in Fig. 4, FL4BCI-SNN demonstrates WQoES improvements of 2.12 and 2.11 times compared to the baseline methods, significantly outperforming both FL4BCI-CNN and FL4BCI-LSTM. This result confirms the superior efficacy of our proposed FL4BCI-SNN model in enhancing WSP. Note that although the constructed WSP metric does not currently integrate with other indicators that influence QoE and QoS, such as resource allocation schemes and wireless communication quality, it is designed to be adaptable and can easily be combined with these factors.
V Challenges and Future Research Directions
In this section, we discuss the practical deployment challenges faced by Bio-HDT from two perspectives: brain signal processing and SNN-driven FL. Additionally, we explore future research directions based on the aforementioned analysis.
V-A Challenges on Processing Brain Signals
The primary challenge in deploying the Bio-HDT system lies in the accurate and real-time processing of brain signals. Given the system’s reliance on continuous monitoring of cognitive, emotional, and behavioral states, the quality of signal extraction and interpretation is paramount. Brain signals, often noisy and prone to external interference, exhibit significant variability between users. Developing robust signal processing techniques that effectively filter out noise while retaining critical features is essential. Additionally, personalized algorithms may be necessary to account for individual differences, as a generic, one-size-fits-all approach may not be effective across diverse user bases.
Another critical factor is ensuring real-time synchronization between brain activity and HDTs within the future human-centric Metaverse. This synchronization is crucial for maintaining an immersive and responsive user experience. With anticipated advancements in B5G/6G systems, the Bio-HDT framework is prone to overcome existing latency challenges by enabling near-instantaneous data transmission. The introduction of 6G could transform Bio-HDT interactions, with ultra-low latency, edge computing, and enhanced data rates combining to ensure seamless and immersive experiences in the Metaverse. These technologies will allow for brain signals to be processed and responded to almost instantly, significantly enhancing user interaction.
V-B Challenges in Implementing SNN-driven FL
One of the primary challenges in deploying SNN-driven FL within the Bio-HDT system lies in the training complexity of SNNs. Unlike traditional ANNs, which process continuous signals, SNNs rely on discrete spikes of electrical activity to transmit information. This discrete, event-driven nature introduces unique difficulties during training. Conventional optimization techniques, such as gradient-based methods (i.e., backpropagation), are less effective for SNNs due to the discontinuous nature of the spike function. Although several approaches, such as surrogate gradient techniques [13] and ANN-to-SNN conversion [12], have been proposed to address these challenges, further exploration is essential to fully unlock the potential of SNNs in real-world applications.
Another significant challenge is the non-independent and identically distributed nature of brain signal data, primarily caused by neurodiversity among users. Devices such as BCI-driven HMDs process brain signals that vary widely due to each individual’s unique neural architecture and cognitive patterns. This variability leads to significant differences in the data across devices, complicating the aggregation and training of a coherent global model in FL. One potential solution is to design cross-learning algorithms that enable efficient transfer learning, allowing the model to adapt to diverse brain signals and improve generalization across users with varying neural characteristics.
VI Conclusion
In this work, we have introduced the current development of HDTs and highlighted the gap between existing HDT frameworks and the evolving requirements of the human-centric Metaverse. To bridge this gap, we have proposed a two-layer bio-inspired scheme. First, we introduce BCI as the first-layer of bio-inspired to create the Bio-HDT framework. This framework is able to capture users’ brain signals, which contain underlying information about humans’ multifaceted data. This capability provides richer data and a holistic view for constructing personalized HDTs, while simultaneously reducing the dependency on the diverse data collection devices. To accurately understand user brain activities, we have proposed a neuromorphic computing learning model based on SNN as the second-layer of bio-inspired. This model not only improves the accuracy of brain signals identification but also reduces energy consumption by mimicking the brain’s natural processing methods, thus enhancing the system performance and then enabling the QoE perceived by the users and QoS provided by MSPs. We have then proposed the SNN-driven FL to enhance privacy conservation. Moreover, we have conducted a case study to validate the efficacy of our proposed scheme. Finally, we have discussed some challenges that are worthy of further research. We believe that this work provides valuable insights into the application of bio-inspired techniques in developing personalized Metaverse services, ultimately advancing the construction of a human-centric next-generation Internet.